# XRISM Spectroscopy of the Stellar-mass Black Hole GRS 1915++105

### v1

@@ 1 Introduction
GRS 1915++105 is a stellar-mass black hole that acts as a bright, local laboratory for understanding accretion physics in active galactic nuclei (AGN). It is sometimes referred to as a “microquasar” for the variety of continuous and discrete radio jet knots that it launches (for a review, see Fender & Belloni 2004). Recently, the impact of these jets on the local interstellar medium has been realized (Motta et al., 2025), potentially revealing scaled versions of the jet cavities that massive black holes inflate in diffuse cluster gas (see, e.g., Fabian et al. 2003). In soft, sub-Eddington, disk-dominated states, GRS 1915++105 launches disk winds with velocities as high as $v=0.03c$ (Miller et al., 2016), similar to “warm absorbers” and “ultra-fast outflows” in Seyferts and quasars (for a review, see Gallo et al. 2023; also see, e.g., Tombesi et al. 2010, Xiang et al. 2025).
@@
The other key properties of GRS 1915++105 are known well, but are continually refined owing to its importance. Radio parallax measurements now place the source at a distance of $d=9.4\pm 0.6$ kpc (Reid & Miller-Jones, 2023). When coupled with prior radial velocity studies, this distance gives a black hole mass of $M=11\pm 2~M_{\odot}$, an inclination of $\theta=64\pm 4$ degrees, and jet velocities between $0.68\leq v/c\leq 0.91$ (Reid & Miller-Jones 2023; also see Reid et al. 2014). Initial fits to X-ray spectra obtained with NuSTAR measured a black hole spin parameter of $a=0.98\pm 0.01$ (where $a=cJ/GM^{2}$, and $J$ is the angular momentum) and an inclination of $\theta=71\pm 1$ degrees (Miller et al., 2013). New fits to multiple spectra with improved models measure a black hole spin parameter of $a=0.98\pm 0.02$ and an inclination of $i=60\pm 8$ degrees (Draghis et al., 2024). It is particularly notable that the fully independent reflection modeling and radio parallax measurements yield commensurate inclination values. Using older values for the black hole mass, distance, and inclination, fits to the X-ray continuum from the disk measure black hole spin parameters of $a\geq 0.98$ (McClintock et al., 2006) and $a\simeq 0.7$ (Middleton et al., 2006).
@@ 3.2 Oscillations in the Fe XXVI RRC?
RRCs arise through the recombination of hot electrons onto atoms. The energy of the emitted photon is the energy of the captured electron, plus the (negative) ionization threshold of the atom. A smooth distribution of electron energies therefore gives rise to a smooth flux decline above the threshold energy. However, a series of eight or more features – spaced by a regular interval of 30 eV – are evident in the spectrum of GRS 1915+++105 (see Figure 5). If the oscillations are real, they are potentially consistent with recombination from electrons in Landau levels in a highly magnetized gas (e.g., in a corona above the disk). However, this explanation is not unique.
@@ 3.3 Results
The results of fits to the Resolve spectrum of GRS 1915++105 with the (non-magnetic) model described above are detailed in Table 1, and shown in Figures 1–5. The model achieves a Cash statistic of $C=3891$ for $\nu=2933$ degrees of freedom. While this does not represent a formally acceptable fit, it is clear in Figures 1–5 that all of the strong emission lines are fit well, that key line ratios are largely reproduced, the evident red-shifts are accurately modeled, and that the predicted RRCs match the data.
@@
Within SPEX, the disk component temperature is measured at the radius of peak emissivity, which exceeds the temperature at the innermost stable circular orbit (ISCO) by a factor of approximately $\simeq 2$. A value of $kT\simeq$1.7-1.8 keV is likely more appropriate for comparisons to XSPEC models (Arnaud, 1996), which measure temperatures at the innermost stable circular orbit (ISCO). Even this temperature is likely higher than the emitted temperature, owing to scattering in gas along the line of sight. In a disk-dominated, unobscured state of GRS 1915++105, Chandra measured an inner disk temperature of $kT=1.6$ keV, and a power-law index of $\Gamma=2.8$ (Miller et al., 2016). The broad similarity of these values to those measured with Resolve – despite very different obscuration profiles – suggests that Resolve observed GRS 1915++105 in an ordinary, sub-Eddington, “thermal-dominant” or “high soft” state, rather than a super- or hyper-Eddington state dominated by a geometrically thick inner disk. (For a review of black hole states, see, e.g., Remillard & McClintock 2006.)
@@
Using a simple $T\propto r^{-3/4}$ scaling and the inferred inner disk temperature, the accretion disk would have a temperature of $kT=0.002$ keV at $r\simeq 1.2\times 10^{5}~GM/c^{2}$. The obscuring gas likely has something to do with the accretion disk itself, since the column is optically thick; however, at least some of the gas might instead be attributed to a warm disk atmosphere above a disk at even larger radii. It is notable that Keplerian orbital velocity at this radius is $\sigma=840~{\rm km}~{\rm s}^{-1}$, close to the rms value of $\sigma=1000~{\rm km}~{\rm s}^{-1}$ that matches the data well. In short, the nature of the obscuration indicates that the outer disk now occults the central engine in GRS 1915++105.
These radii are not compatible with the width of the neutral Fe Kα line. The observed blending of the two line components and the absence of an observable line shift require a combination of a smaller radius and a low inclination. In broad terms, the neutral Fe Kα line originates 1–2 orders of magnitude closer to the central engine than the obscuring gas does (see Table 1). The low inclination of the neutral Fe Kα line, $i=5^{+4}_{-2}$ degrees, is very different from that of the binary system ($i=64\pm 4$ degrees, Reid & Miller-Jones 2023), the innermost disk ($i=60\pm 8$ degrees, Draghis et al. 2024), and the jet ($i=66\pm 2$ degrees, Fender et al. 1999). Indeed, if the disk is flat, we should not see particular radii evident as a sharp line. These conditions needed to produce the neutral Fe Kα line are broadly consistent with a warp that first departs from a flat disk at $r\simeq 10^{3-4}~GM/c^{2}$, bringing the outer disk into the line of sight by $r\simeq 10^{5-6}~GM/c^{2}$.
@@
Thermal winds can be launched outside of the Compton radius, given by ${\rm R}_{\rm C}=1\times 10^{10}\times({\rm M}_{BH}/{\rm M}_{\odot})\times{\rm T}_{\rm C,8}^{-1}$ (where ${\rm T}_{\rm C,8}$ is the Compton temperature in units of $10^{8}$ K; Begelman et al. 1983). Approximating the Compton temperature as the disk temperature, ${\rm R}_{\rm C}=1.6\times 10^{5}~GM/c^{2}$, inside the radius at which the the photoionized lines are produced. It is therefore plausible that the narrow lines trace a thermal wind. Indeed, the large inferred distance from the central engine, high ionizations, and low velocities are all broadly consistent with numerical simulations of thermal winds (see, e.g., Higginbottom et al. 2017), but the inconsistent shifts may again point to potential axial asymmetry. The broad line that underlies the narrow Fe XXV He-$\alpha$ complex may not be part of a thermal wind, given an origin at $r\simeq 3\times 10^{3}~GM/c^{2}$.
@@ 3.4 The Fe XXVI RRC
The best-fit magnetic field is measured to be ${\rm B}=2.57\pm 0.06\times 10^{9}$ Gauss. The phase of the observed oscillations is measured to be $\phi=0.13^{+0.08}_{-0.03}$, and the amplitude is measured to be $A=1.9\pm 0.4$. This field is extremely high, commensurate with the surface magnetic field inferred in accreting millisecond X-ray pulsars (Patruno & Watts, 2021). The putative field would have to exist in a hot, electron-dominated corona that may differ from the underlying accretion disk. However, assuming that these features are also produced in the outermost accretion disk, the implied magnetic field is several orders of magnitude greater than equipartition between magnetic pressure and gas pressure (Shakura & Sunyaev 1973; also see Miller et al. 2016). In short, a field this high would easily disrupt the disk. It is not clear how an equipartition field that might arise in the disk could be amplified to such a degree in a blanketing atmosphere or corona.
Moreover, the apparent oscillations are quite sharp. Whatever the exact orientation of the corona and field, separations of 30 eV would be completely blurred out at radii less than $r\leq 1.1\times 10^{5}~GM/c^{2}$ and so much larger radii – $r\simeq 1\times 10^{6}~GM/c^{2}$ – are likely. At this radius, however, the corona would be unbound, and could represent an ionized outflow. The distinct velocities are qualitatively similar to the wind observed in PDS 456, and so are the velocities (Xrism Collaboration et al., 2025). The lines cannot be dominated by H-like Fe XXVI, because the 18 eV spacing of the line components is incompatible with the data. Rather, intermediate charge states would be required, and velocities up to $v=0.34c$. If the wind is launched verticaly, the true velocities would even higher.
@@ 4 Discussion
We have analyzed the first XRISM/Resolve spectrum of the stellar-mass black hole GRS 1915++105. The source was captured in its “obscured” state. Our models suggest that the central engine remains in a disk–dominated, sub-Eddington accretion mode, but is largely occulted by the outer accretion disk. This situation is qualitatively similar to highly obscured and Compton-thick AGN, wherein the central engine is obscured by a molecular torus at the outer edge of the accretion flow. In this section, we discuss the nature of the “obscured” state in more detail, considering multi-wavelength observations and analogous sources. We also discuss the shortcomings of our model, alternative emission mechanisms that may contribute to the strong emission line spectrum, and comment on apparent structure in the Fe XXVI RRC.
@@ 4.1 The nature of the obscured state
A super-Eddington accretion rate could cause the inner accretion disk to assume a “funnel” geometry (King et al., 2001), blocking some of the ionizing flux and potentially accounting for some aspects of the “obscured” state. This configuration remained a possibility based on Chandra gratings spectra obtained in the “obscured” state, which did not require an accretion disk. The much higher sensitivity achieved even in this short Resolve spectrum does require a disk, but the disk temperature is typical of its standard, sub-Eddington “thermal-dominant” state (historically called the “high/soft” state; for a review of states, see Remillard & McClintock 2006).
@@ 4.4 Open questions
4.5 Explaining the Fe XXVI RRC
JMM acknowledges helpful conversations with Ehud Behar, James Miller-Jones, Jelle Kaastra, Caroline Kilbourne, Frits Paerels, Daniel Proga, Luis Rodriguez, Daniele Rogantini, and Sergei Trushkin. JMM thanks Jelle de Plaa for sharing advanced versions of SPEX, and thanks Savannah Ware, Elissa Slotkin, Sydney Hess, and Mark Kelly for their efforts to keep NASA astrophysics vibrant, including support for XRISM.

### v2

@@ 1 Introduction
GRS 1915++105 is a stellar-mass black hole that acts as a bright, local laboratory for understanding accretion physics in active galactic nuclei (AGN). It is sometimes referred to as a “microquasar” for its occasional, apparently superluminal jet velocity, and the variety of continuous and discrete radio jets that it launches (e.g., Mirabel & Rodríguez 1994; for a review, see Fender & Belloni 2004). Recently, the impact of these jets on the local interstellar medium has been realized (Motta et al., 2025), potentially revealing scaled versions of the jet cavities that massive black holes inflate in diffuse cluster gas (see, e.g., Fabian et al. 2003).
GRS 1915++105 is also known for a broad array of at least 12 variability patterns that may reflect particular configurations of the disk, corona, and jet (Belloni et al., 2000). The character of disk winds in GRS 1915++105 appears to vary with spectral state and the degree of radio jet activity (see, e.g., Neilsen & Lee 2009, but some wind activity may persist in all states (Neilsen et al., 2018). However, in soft, sub-Eddington, disk-dominated states, GRS 1915++105 launches disk winds with velocities as high as $v=0.03c$ (Miller et al., 2016), similar to “ultra-fast outflows” in Seyferts and quasars (for a review, see Gallo et al. 2023; also see, e.g., Tombesi et al. 2010, Xiang et al. 2025).
@@
The other key properties of GRS 1915++105 are known well, but are continually refined owing to its importance. Radio parallax measurements now place the source at a distance of $d=9.4\pm 0.6$ kpc (Reid & Miller-Jones, 2023). When coupled with prior radial velocity studies, this distance gives a black hole mass of $M=11\pm 2~M_{\odot}$, an inclination of $\theta=64\pm 4$ degrees, and jet velocities between $0.68\leq v/c\leq 0.91$ (Reid & Miller-Jones 2023; also see Reid et al. 2014). Owing to the extreme nature its jets and the possibility that they are powered by the spin of the black hole (e.g., Blandford & Znajek 1977), considerable effort has also been devoted to measuring this aspect of GRS 1915++105. The higher radiative efficiency of accretion onto a highly spinning black hole may be important to the obscured state of GRS 1915++105, as the enhanced radiative output may enable thickening in the outer disk or the formation a radiation-induced warp (e.g., Pringle 1996).
Initial fits to X-ray spectra obtained with NuSTAR measured a black hole spin parameter of $a=0.98\pm 0.01$ (where $a=cJ/GM^{2}$, and $J$ is the angular momentum) and an inclination of $\theta=71\pm 1$ degrees (Miller et al., 2013). New fits to multiple spectra with improved models measure a black hole spin parameter of $a=0.98\pm 0.02$ and an inclination of $i=60\pm 8$ degrees (Draghis et al., 2024). It is particularly notable that the fully independent reflection modeling and radio parallax measurements yield commensurate inclination values. Using older values for the black hole mass, distance, and inclination, fits to the X-ray continuum from the disk measure black hole spin parameters of $a\geq 0.98$ (McClintock et al., 2006) and $a\simeq 0.7$ (Middleton et al., 2006). The latter of these matches the value of $a=0.71\pm 0.03$ derived from associating 67 Hz quasi-periodic oscillations (QPOs) observed in GRS 1915++105 with nodal precession (Motta & Belloni, 2024). This value is twice higher than spins derived in other stellar-mass black holes using the same method.
@@ 3.2 Oscillations in the Fe XXVI RRC?
RRCs arise through the recombination of hot electrons onto atoms. The energy of the emitted photon is the energy of the captured electron, plus the (negative) ionization threshold of the atom. A smooth distribution of electron energies therefore gives rise to a smooth flux decline above the threshold energy. However, a series of eight or more features – spaced by a regular interval of 30 eV – are evident in the spectrum of GRS 1915++105 (see Figure 5). If the oscillations are real, they are potentially consistent with recombination from electrons in Landau levels in a highly magnetized gas (e.g., in a corona above the disk). However, this explanation is not unique.
@@ 3.3 Results
The results of fits to the Resolve spectrum of GRS 1915++105 with the (non-magnetic) model described above are detailed in Table 1, and shown in Figures 1–5. The model achieves a Cash statistic of $C=3891$ for $\nu=2933$ degrees of freedom. While this does not represent a formally acceptable fit, it is clear in Figures 1–5 that all of the strong emission lines are fit well, that key line ratios are largely reproduced, the evident red-shifts are accurately modeled, the predicted RRCs match the data, and the continuum is modeled well.
@@
Within SPEX, the disk component temperature is measured at the radius of peak emissivity, which exceeds the temperature at the innermost stable circular orbit (ISCO) by a factor of approximately $\simeq 2$. A value of $kT\simeq$1.7-1.8 keV is likely more appropriate for comparisons to XSPEC models (Arnaud, 1996), which measure temperatures at the ISCO. Even this temperature is likely higher than the emitted temperature, owing to scattering in gas along the line of sight. In a disk-dominated, unobscured state of GRS 1915++105, Chandra measured an inner disk temperature of $kT=1.6$ keV, and a power-law index of $\Gamma=2.8$ (Miller et al., 2016). The broad similarity of these values to those measured with Resolve – despite very different obscuration profiles – suggests that Resolve observed GRS 1915++105 in an ordinary, sub-Eddington, “thermal-dominant” or “high soft” state, rather than a super- or hyper-Eddington state dominated by a geometrically thick inner disk. (For a review of black hole states, see, e.g., Remillard & McClintock 2006.)
@@
Using a simple $T\propto r^{-3/4}$ scaling and the inferred inner disk temperature, the accretion disk would have a temperature of $kT=0.002$ keV at $r\simeq 1.2\times 10^{5}~GM/c^{2}$. The obscuring gas likely has something to do with the accretion disk itself, since the column is optically thick; however, at least some of the gas might instead be attributed to a warm disk atmosphere above a disk at even larger radii. It is notable that Keplerian orbital velocity at this radius is $\sigma=840~{\rm km}~{\rm s}^{-1}$, close to the rms value of $\sigma=1000~{\rm km}~{\rm s}^{-1}$ that matches the data well. In short, the nature of the obscuration is consistent with the outer disk occulting the central engine in GRS 1915++105.
These radii are not compatible with the width of the neutral Fe Kα line. The observed blending of the two line components and the absence of an observable line shift require a combination of a smaller radius and a low inclination. In broad terms, the neutral Fe Kα line originates 1–2 orders of magnitude closer to the central engine than the obscuring gas does (see Table 1). The low inclination of the neutral Fe Kα line, $i=5^{+4}_{-2}$ degrees, is very different from that of the binary system ($i=64\pm 4$ degrees, Reid & Miller-Jones 2023), the innermost disk ($i=60\pm 8$ degrees, Draghis et al. 2024), and the jet ($i=66\pm 2$ degrees, Fender et al. 1999). Indeed, if the disk is flat and the central engine can be treated as a point source, each annulus should contribute progressively less emission, and we should expect to see a broad, indistinct line. Sharp lines from specific radii – e.g., from the broad line region or inner face of the torus in AGN (Xrism Collaboration et al., 2024) – require a geometry that adds more emitting area per radius than a truly flat disk can produce. These conditions needed to produce the neutral Fe Kα line are broadly consistent with a warp that first departs from a flat disk at $r\simeq 10^{3-4}~GM/c^{2}$, bringing the outer disk into the line of sight by $r\simeq 10^{5-6}~GM/c^{2}$.
@@
Thermal winds can be launched outside of the Compton radius, given by ${\rm R}_{\rm C}=1\times 10^{10}\times({\rm M}_{BH}/{\rm M}_{\odot})\times{\rm T}_{\rm C,8}^{-1}$ (where ${\rm T}_{\rm C,8}$ is the Compton temperature in units of $10^{8}$ K; Begelman et al. 1983). Approximating the Compton temperature as the disk temperature, ${\rm R}_{\rm C}=1.6\times 10^{5}~GM/c^{2}$, inside the radius at which the photoionized lines are produced. It is therefore plausible that the narrow lines trace a thermal wind. Indeed, the large inferred distance from the central engine, high ionizations, and low velocities are all broadly consistent with numerical simulations of thermal winds (see, e.g., Higginbottom et al. 2017), but the inconsistent shifts may again point to potential axial asymmetry. The broad line that underlies the narrow Fe XXV He-$\alpha$ complex may not be part of a thermal wind, given an origin at $r\simeq 3\times 10^{3}~GM/c^{2}$.
@@ 3.4 The Fe XXVI RRC
The best-fit magnetic field is measured to be ${\rm B}=2.57\pm 0.06\times 10^{9}$ Gauss. The phase of the observed oscillations is measured to be $\phi=0.13^{+0.08}_{-0.03}$, and the amplitude is measured to be $A=1.9\pm 0.4$. This field is extremely high, commensurate with the surface magnetic field inferred in accreting millisecond X-ray pulsars (Patruno & Watts, 2021). The putative field would have to exist in a hot, electron-dominated corona that may differ from the underlying accretion disk. However, assuming that these features are also produced in the outermost accretion disk, the implied magnetic field is several orders of magnitude greater than equipartition between magnetic pressure and gas pressure (Shakura & Sunyaev 1973; also see Miller et al. 2016). A field this high would easily disrupt the disk, and it is not clear how an equipartition field that might arise in the disk could be amplified to such a degree in a blanketing atmosphere or corona; in summary, the features may not be due to magnetic effects (also see Section 4.5, below).
Moreover, the apparent oscillations are quite sharp. Whatever the exact nature of the gas, separations of 30 eV would be completely blurred out at radii less than $r\leq 1.1\times 10^{5}~GM/c^{2}$ and so much larger radii – $r\simeq 1\times 10^{6}~GM/c^{2}$ – are likely. At this radius, however, the corona would be unbound, and could represent an ionized outflow. The distinct velocities are qualitatively similar to the wind observed in PDS 456, and so are the velocities (Xrism Collaboration et al., 2025). The lines cannot be dominated by H-like Fe XXVI, because the 18 eV spacing of the line components is incompatible with the data. Rather, intermediate charge states would be required, and velocities up to $v=0.34c$. If the wind is launched verticaly, the true velocities would be even higher.
@@ 4 Discussion
We have analyzed the first XRISM/Resolve spectrum of the stellar-mass black hole GRS 1915++105. The source was captured in its “obscured” state. Our models suggest that the central engine remains in a disk–dominated, sub-Eddington accretion mode, but is largely occulted by the outer accretion disk. This situation is qualitatively similar to highly obscured and Compton-thick AGN, wherein the central engine is obscured by the torus at the outer edge of the accretion flow. However, there are important differences, including that the torus is cold enough to retain dust and molecules, whereas the outer disk in X-ray binaries is not. In this section, we discuss the nature of the “obscured” state in more detail, considering multi-wavelength observations and analogous sources. We also discuss the shortcomings of our model, alternative emission mechanisms that may contribute to the strong emission line spectrum, and comment on apparent structure in the Fe XXVI RRC.
@@ 4.1 The nature of the obscured state
A super-Eddington accretion rate could cause the inner accretion disk to assume a “funnel” geometry (King et al., 2001), blocking some of the ionizing flux and potentially accounting for some aspects of the “obscured” state. This configuration remained a possibility based on Chandra gratings spectra obtained in the “obscured” state, which did not require an accretion disk. The much higher sensitivity achieved even in this short Resolve spectrum does require a disk, but the disk temperature is typical of bright, sub-Eddington soft states (e.g., Homan & Belloni 2005, also see Remillard & McClintock 2006).
@@ 4.5 Open questions
5 Summary and Conclusions
The results of our analysis of the first Resolve spectrum of GRS 1915++105 in its “obscured state” can be summarized as follows:
$\bullet$ The central engine in GRS 1915++105 is likely obscured by warm, partially covering, Compton-thick gas.
$\bullet$ After accounting for this obscuration, the data are consistent with a standard, sub-Eddington, inner accretion flow that is dominated by a hot accretion disk.
$\bullet$ The high-resolution calorimeter spectrum is dominated by strong, narrow emission lines from He-like and H-like charge states, consistent with photoionization in the outer accretion disk.
$\bullet$ It is therefore likely that the outer disk obscures the central engine. This could be due to axially symmetric thickening of the outer disk in response to a higher central luminosity than we have inferred, or due to warping and precession, similar to super-orbital periods and low-flux phases in other wide binary systems. The latter explanation is potentially consistent with observed changes in the position angle and inclination angle of the radio jet in GRS 1915++105 (Rodríguez & Mirabel, 2025).
$\bullet$ The observed emission lines likely probe the atmosphere of the outer accretion disk, or the base of a slow wind (but note that these may not be entirely distinct). The measured abundances suggest enhancements in Cr and Mn, and are broadly consistent with models wherein such elements are deposited into the companion star by the black hole creation event (Nomoto et al. 2013, Keshet et al. 2025).
The authors thank the anonymous referee for helpful comments that improved this manuscript. JMM acknowledges helpful conversations with Ehud Behar, James Miller-Jones, Jelle Kaastra, Caroline Kilbourne, Frits Paerels, Daniel Proga, Luis Rodriguez, and Sergei Trushkin. JMM thanks Jelle de Plaa for sharing advanced versions of SPEX, and thanks Savannah Ware, Elissa Slotkin, Sydney Hess, and Mark Kelly for their efforts to keep NASA astrophysics vibrant, including support for XRISM.